OpenStudy (anonymous):

A sailor pulls a boat horizontally alongside a dock for 10.0 m with a rope angled at 40.0°. He uses a force of 100.0 N. How much work did he do on the boat? A. 766 J B. 643 J C. 1.00 * 10^3 J D. 4.00 * 10^4 J

OpenStudy (michele_laino):

here is the situation described in your problem: 

OpenStudy (anonymous):

yes:)

OpenStudy (michele_laino):

now, we have to apply this formula:

OpenStudy (anonymous):

which formula?

OpenStudy (michele_laino):

\[W = Fd\cos \left( {40} \right) = 100 \times 10 \times \cos 40\]

OpenStudy (anonymous):

766.0444 so about 766 J is our answer?

OpenStudy (michele_laino):

yes! That's right!

OpenStudy (anonymous):

yay!! :D

OpenStudy (michele_laino):

:D